The difference of two numbers is 18. Find the numbers such that their product is a minimum.
Answer by stanbon(57219)

The difference of two numbers is 18. Find the numbers such that their product is a minimum.
---------------------
x-y = 18
------------------
product = xy
---
y = x-18
--
Substitute and determine minimum product.
----
P = x(x-18)
P = x^2-18x
---
Minimum occurs when x = -b/(2a) = 18/(2) = 9
y = x-18
y = 9-18 = -9
-----
Minimum product is xy = 9*-9 = -81
=====================================
Cheers,
Stan H.
==========